# A Priori Assessment of Rotational Invariance in Multiscale CNN-Based Subgrid-Scale Model for Wall-Bounded Turbulent Flows

### v1

@@ I Introduction
Large-eddy simulation (LES) is a numerical approach for accurately capturing the dynamics of turbulent flows in both engineering and science applications. In contrast to direct numerical simulation (DNS), which resolves all scales of motion, LES resolves large-scale or grid-scale (GS) dynamics and represents the small-scale or subgrid-scale (SGS) structures using the SGS model. Conventional SGS models, especially eddy-viscosity models (EVM), have the advantages of simplicity and robustness. However, the inherent linearity of an EVM may limit its accuracy, especially in capturing anisotropic turbulencespalartphilosophies2015. Efforts made to address this limitation include the scale-similarity (SS) model bardinaimproved1980 and nonlinear model or Clark model clarkevaluation1979. In the SS model, the SGS structures are assumed to be similar across neighboring scales, meaning that the flow structures observed at different filtering levels are similar to one another. This model employs a test filter to approximate the SGS stresses ($\tau_{ij}$) through the modified Leonard tensorbardinaimproved1980. The nonlinear model is based on SGS stresses explicitly derived from a Taylor-series expansion of the filtering operation. Both models reconstruct the SGS stresses directly without invoking the eddy-viscosity assumption. Although these models are well-correlated with DNS data, they tend to lack sufficient dissipationkajishima2017; sagaut2006. Advancements such as the mixed modelBardina1983, which combines the SS model with Smagorinsky modelsmag1963, and the dynamic Smagorinsky model (DSM)germano, have been proposed. The mixed model leverages the local structural accuracy of the SS model and introduces a dissipation term from the Smagorinsky model. This model remains constrained by the assumptions underlying each of the SS and Smagorinsky componentskajishima2017; sagaut2006. Germano et al.germano proposed the SGS model for adapting the Smagorinsky model by dynamically adjusting the Smagorinsky constant at each point in space and time to improve the adaptability of the model to the local flow structure. The computation of this constant aims to avoid a priori estimation of the Smagorinsky model by using the Germano identity with both the grid filter and an additional test filtergermano. Nevertheless, this model suffers from instability owing to negative coefficients, and numerical averaging is required to achieve stabilitysagaut2006; kajishima2017. However, this averaging offsets the main advantage of the DSM to determine the local and dynamic Smagorinsky constants directly from the resolved flow fieldkajishima2017.
@@
Owing to the challenges faced by conventional SGS models, data-driven strategies have recently gained significant attention as an alternative approach for reconstructing SGS models. Deep neural networks (DNN), trained on spatially filtered DNS (fDNS) datasets, have demonstrated the ability to learn mappings from resolved flow variables to the unclosed $\tau_{ij}$. In contrast to conventional SGS models, data-driven approaches aim to inductively extract small-scale SGS features without relying on artificial assumptions. To the best of the authors’ knowledge, the earliest application of a data-driven SGS model was proposed by Sarghini et al. Sarghini2003, who estimated the eddy-viscosity coefficient using a multilayer perceptron (MLP) instead of directly calculating $\tau_{ij}$ in the momentum equation. Their results were comparable to those of the conventional Bardina model in terms of the turbulence statisticsSarghini2003. Subsequently, in contrast to the approach of Sarghini et al. Sarghini2003, who employed a shallow DNN, Pal pal2019deeplearningparameterizationsubgrid developed an MLP-based SGS model with a deeper DNN architecture to reconstruct the eddy-viscosity coefficient. This deeper architecture allowed the model to represent more complex mappings between the resolved flow variables and $\tau_{ij}$, thereby improving its ability to capture the turbulence features. As that of result, their findings demonstrated improved performance over the DSM in channel-flow simulations, with the turbulence statistics demonstrating a closer agreement with the DNS datapal2019deeplearningparameterizationsubgrid. Yet, the effectiveness of these approaches is often attributed to their dependence on the alignment with the rate-of-strain tensor $\bar{D}_{ij}$, as the accuracy hinges on the eddy-viscosity formulation, which is a form of artificial assumptionPope2000. Here, the overbar $\overline{(\cdot)}$ denotes the filtering operation.
@@
Rather than inferring the eddy-viscosity coefficients, researchers have adopted a direct-closure strategy by predicting $\tau_{ij}$ directly from the resolved turbulent fields (e.g., Refs. beck2019; kurz2022; gamahara2017; zhou2019; liu2022; xie2020; park2021; guan2022; jalaali2025). These approaches envision functions that are not constrained by artificial assumptions, such as an EVM, thus fully leveraging the benefits of data-driven models. Beck et al. beck2019 suggested that the data-driven SGS model is feasible as long as the direct-closure terms are obtained from a previously conducted DNS. Accordingly, Kurz &\& Beck kurz2022 employed an MLP-based SGS model and showed a good agreement with the fDNS for isotropic turbulence. A similar approach has been adopted by several researchers, such as Gamahara and Hattori gamahara2017, Zhou et al. zhou2019, and Liu et al. liu2022, who predicted $\tau_{ij}$ using resolved flow variables as the input to their MLP-based SGS model. Their results demonstrated that this direct-closure strategy could reproduce the SGS stress distribution in a priori test and turbulence statistics in a posteriori test with a close agreement with the result of the fDNS.
@@
In subsequent studies, instead of applying an MLP algorithm in which both input and output are defined at single-point locations, researchers such as Xie et al. xie2020 and Park &\& Choi park2021 have explored multipoint inputs to improve the predictive accuracy. This approach enables the MLP algorithm to capture turbulent interactions by accounting for non-local input features. Xie et al. xie2020 used multipoint data of the velocity gradient $\frac{\partial\bar{u}_{i}}{\partial x_{j}}$ as the input and achieved accurate a priori predictions, along with improvements in a posteriori tests when compared with the results of the conventional DSM for isotropic turbulence. Park &\& Choi park2021 evaluated both single- and multipoint inputs in a channel flow case. They found that multipoint inputs improved the $\tau_{ij}$ prediction accuracy and achieved a good agreement with the fDNS result in a posteriori test. However, it is worth noting that MLP-based algorithms with multiple inputs often impose a substantial computational burden, making them less effective for large-scale datasetsliu2022.
@@
A more efficient approach to extract multipoint spatial information is through the use of convolutional neural networks (CNN), which capture the spatial correlations via convolutional kernels such that both the memory and computational cost can be reduced, making them more suitable for flow-field datamorimoto2021; liu2022. Given these advantages, CNN-based SGS models have been employed in previous studies, such as those of Beck et al. beck2019 and Guan et al. guan2022, for homogeneous isotropic turbulence. Their results demonstrated accurate predictions of $\tau_{ij}$ and showed that the CNN-based models outperformed both static and dynamic Smagorinsky modelsbeck2019; guan2022 and the MLP-based SGS models in a posteriori test. Liu et al. liu2022 employed a CNN-based data-driven SGS model for a turbulent channel flow. They also compared the CNN-based and MLP-based SGS models, showing that the CNN-based model provided more accurate predictions of $\tau_{ij}$. In a posteriori tests, the CNN-based model outperformed the DSM in terms of turbulence statistics. This improvement underscores the advantage of using CNN to capture spatial and turbulent interactions across the neighboring grid points. Furthermore, it is crucial to represent the multiscale characteristics of turbulent flow dynamicsxie2020; jalaali2025. Xie et al.xie2020 suggested that data-driven SGS models should capture the multiscale characteristic and energy cascade of turbulence through appropriately designed multiple-point input features. To this end, Jalaali and Okabayashi jalaali2025 introduced a multiscale CNN-based SGS model designed to capture the vortex interactions across scales, which are essential for energy and momentum transfer. Their model achieved more accurate predictions of $\tau_{ij}$ and improved the turbulent statistics in a posteriori tests when compared with those of both the conventional Smagorinsky model and other CNN-based data-driven SGS models. Moreover, it reproduced the inverse energy cascade and maintained numerical stability in a posteriori testsjalaali2025. Such improvements are likely attributable to the ability of the model to capture interscale vortex interactions, which are not explicitly represented in conventional CNN-based approaches.
@@
Despite these promising developments, for any data-driven SGS model to be physically valid, it must adhere to fundamental invariance principles inherent to turbulence modeling. In modeling terms, the turbulence model must be consistent and frame-indifferent, or satisfy the principle of material objectivity. Fulfilling this requirement has been treated as a hard constraint not only for conventional SGS modelskajishima2017 but also the data-driven SGS models Spalart2023; duraisamy2021. As discussed by OberlackOberlack1997, classical SGS models often fail to meet this requirement. In particular, models that explicitly include the rate-of-rotation tensor frequently violate the requirement of frame indifference. To address this, considerable effort has been made to incorporate invariance into the SGS formulations. For example, OberlackOberlack1997 noted that the DSM proposed by Germano et al.germano satisfied this principle by introducing invariant variables and test-filter quantities that preserved the scaling invariance. Although the model structures considerably differ from those of the conventional SGS models, data-driven SGS models must also adhere to material objectivity, ensuring physically consistent behavior under coordinate rotations. Accordingly, data-driven SGS models are expected to respect the following transformation property:
@@
Likewise, the efforts to satisfy the material objectivity have been undertaken within the data-driven turbulence modeling framework. In this regard, the generalized expansion of the Reynolds stress tensor by Pope pope1975 has been employed to develop invariant closures for Reynolds stresses. This approach is based on the assumption that the stresses depend on the mean velocity gradient Pope2000. Within the data-driven framework, this model rewrites the expansion of the Reynolds stress tensor by Pope as
@@
where $T^{(n)}$ are known functions of the symmetric and antisymmetric parts of the velocity gradient tensor, and $\lambda_{1},\ldots,\lambda_{5}$ are scalar invariants. The detailed derivation of these 10 tensors and 5 invariants is given in PopePope2000; pope1975. The goal is to determine the corresponding coefficients ($c^{(n)}(\lambda_{1},\ldots,\lambda_{5})$), following which the Reynolds stress tensor can be calculated using Eq. (2). The data-driven approach was initially introduced by Ling et al. Lingmachine2016; LingReynolds2016 as the tensor basis neural network (TBNN), and later extended by other researchers, such as Brener et al. brener2024 in the context of data-driven RANS modeling and Bose &\& Roybose2024 for LES closures. These studies relied on scalar invariants to reconstruct the anisotropic stresses and showed a relatively good agreement with the fDNS data.
@@
However, there are limitations to the TBNN approach. As noted by Cinnella cinnella2024, the TBNN based on the representation by Pope relies on the hypothesis of the local equilibrium of turbulence, an assumption that no longer holds for flows with strong non-equilibrium. To some extent, the polynomial expansion used in the representation by Pope can be extended to better capture the non-equilibrium turbulencepope1975; wu1028. Prakash et al. prakash2022 noted that the size of the minimal tensor set and the corresponding invariant basis increases with the number of prescribed input features. To address this issue, Prakash et al. prakash2022 proposed a data-driven SGS model based on the eigenstructure of the SGS tensor, using invariant variables as inputs. Their results showed a good agreement with the DNS data for isotropic turbulence. Brener et al. brener2024 and Wu et al.wu1028 showed that invariance can also be achieved through data augmentation. By exposing the DNN architecture to multiple orientations of the dataset, the data-driven SGS model was expected to learn representations that are insensitive to frame rotationswu1028. However, data augmentation was computationally expensive because it required duplicating the DNS datasets across multiple transformed coordinate systemsbrener2024. Beyond input design and data augmentation, a critical concern remains that invariance should be preserved not only at the level of the input features but also within the structure and operations of the DNN algorithm itselfJaderberg; prakash2022; bose2024. For example, the study by Jalaali &\& Okabayashijalaali2025 used rotationally invariant inputs, but the architecture of the DNN model itself did not maintain rotational invariance as, a CNN algorithm cannot preserve rotational invariance without the necessary treatmentgoodfellow2016. Because DNN is inherently composed of coordinate-sensitive operations, it is desirable to design architectures or training strategies that are intrinsically invariant.
@@
Motivated by these challenges, we developed a data-driven SGS model that enforces material objectivity and preserves rotational invariance. Here, we adopted the multiscale CNN-based SGS model introduced by Jalaali &\& Okabayashi jalaali2025 (MSC model), which captures the flow features across multiple spatial scales. The contributions of this study can be outlined as follows. First, we modified the MSC model by removing the bias terms and batch normalization (BN) layers, following Bin et al. Bin2022 and Cho et al. Cho2024, allowing the model to operate directly on raw input data and enhancing the invariance properties. Second, we integrated a spatial transformer network (STN) Jaderberg into the architecture, enabling the model to learn the features that remain consistent under coordinate rotations. These design choices were evaluated through a priori test with rotated inputs to examine the rotational invariance. The remainder of this paper is organized as follows. In Section II, we introduce the numerical framework and present the proposed data-driven SGS model, with particular emphasis on the overall modeling strategy. Section III describes the results and discussions. Finally, Section IV concludes the study by summarizing the key findings and discussing possible directions for future developments.
@@ II.1 Problem setting
The residual SGS stress, $\tau_{ij}=\overline{u_{i}u_{j}}-\bar{u}_{i}\bar{u}_{j}$, is unclosed and represents the effect of unresolved scales. Therefore, a closure model is required; while by far the most widely used closure adopts the eddy-viscosity hypothesis that can be modeled using conventional approaches such as the Smagorinsky modelsmag1963. In the present study, we replaced such conventional closures with a data-driven SGS model for $\tau_{ij}$. Fig. 1 outlines the modeling workflow. This study was limited to a priori assessment of the capability of the proposed data-driven SGS model to achieve rotational invariance. In a priori tests, the data-driven SGS model is trained in a supervised manner on fDNS flow fields. By using the input data X{X} and label data $\tau_{ij}$, the data-driven SGS model learns a nonlinear mapping such that $\tau_{ij}^{p}=\mathcal{F}({X};w)$, where $\tau_{ij}^{p}$ is the predicted $\tau_{ij}$. The parameter $w$ denotes the weights of the DNN and is obtained by optimization, given as $w=\arg\min_{w}\left(\mathcal{L}(\tau_{ij},\tau_{ij}^{p})\right)$. $\mathcal{L}$ denotes the loss function.

### v2

# A Priori Assessment of Rotational Invariance in Multiscale Convolutional Neural Network-Based Subgrid-Scale Model for Wall-Bounded Turbulent Flows

@@ I Introduction
Large-eddy simulation (LES) is a numerical approach for accurately capturing the dynamics of turbulent flows in both engineering and science applications. In contrast to direct numerical simulation (DNS), which resolves all scales of motion, LES resolves large-scale or grid-scale (GS) dynamics and represents the small-scale or subgrid-scale (SGS) structures using the SGS model. Conventional SGS models, especially eddy-viscosity models (EVM), have the advantages of simplicity and robustness. However, the inherent linearity of an EVM may limit its accuracy, especially in capturing anisotropic turbulence[41]. Efforts made to address this limitation include the scale-similarity (SS) model [3] and nonlinear model or Clark model [10]. In the SS model, the SGS structures are assumed to be similar across neighboring scales, meaning that the flow structures observed at different filtering levels are similar to one another. This model employs a test filter to approximate the SGS stresses ($\tau_{ij}$) through the modified Leonard tensor[3]. The nonlinear model is based on SGS stresses explicitly derived from a Taylor-series expansion of the filtering operation. Both models reconstruct the SGS stresses directly without invoking the eddy-viscosity assumption. Although these models are well-correlated with DNS data, they tend to lack sufficient dissipation[22, 38]. Advancements such as the mixed model[2], which combines the SS model with Smagorinsky model[40], and the dynamic Smagorinsky model (DSM)[15], have been proposed. The mixed model leverages the local structural accuracy of the SS model and introduces a dissipation term from the Smagorinsky model. This model remains constrained by the assumptions underlying each of the SS and Smagorinsky components[22, 38]. Germano et al.[15] proposed the SGS model for adapting the Smagorinsky model by dynamically adjusting the Smagorinsky constant at each point in space and time to improve the adaptability of the model to the local flow structure. The computation of this constant aims to avoid a priori estimation of the Smagorinsky model by using the Germano identity with both the grid filter and an additional test filter[15]. Nevertheless, this model suffers from instability owing to negative coefficients, and numerical averaging is required to achieve stability[38, 22]. However, this averaging offsets the main advantage of the DSM to determine the local and dynamic Smagorinsky constants directly from the resolved flow field[22].
@@
Owing to the challenges faced by conventional SGS models, data-driven strategies have recently gained significant attention as an alternative approach for reconstructing SGS models. Deep neural networks (DNN), trained on spatially filtered DNS (fDNS) datasets, have demonstrated the ability to learn mappings from resolved flow variables to the unclosed $\tau_{ij}$. In contrast to conventional SGS models, data-driven approaches aim to inductively extract small-scale SGS features without relying on artificial assumptions. To the best of the authors’ knowledge, the earliest application of a data-driven SGS model was proposed by Sarghini et al. [39], who estimated the eddy-viscosity coefficient using a multilayer perceptron (MLP) instead of directly calculating $\tau_{ij}$ in the momentum equation. Their results were comparable to those of the conventional Bardina model in terms of the turbulence statistics[39]. Subsequently, in contrast to the approach of Sarghini et al. [39], who employed a shallow DNN, Pal [33] developed an MLP-based SGS model with a deeper DNN architecture to reconstruct the eddy-viscosity coefficient. This deeper architecture allowed the model to represent more complex mappings between the resolved flow variables and $\tau_{ij}$, thereby improving its ability to capture the turbulence features. As that of result, their findings demonstrated improved performance over the DSM in channel-flow simulations, with the turbulence statistics demonstrating a closer agreement with the DNS data[33]. Yet, the effectiveness of these approaches is often attributed to their dependence on the alignment with the rate-of-strain tensor $\bar{D}_{ij}$, as the accuracy hinges on the eddy-viscosity formulation, which is a form of artificial assumption[36]. Here, the overbar $\overline{(\cdot)}$ denotes the filtering operation.
@@
Rather than inferring the eddy-viscosity coefficients, researchers have adopted a direct-closure strategy by predicting $\tau_{ij}$ directly from the resolved turbulent fields (e.g., Refs. [4, 26, 14, 46, 29, 45, 34, 17, 21]). These approaches envision functions that are not constrained by artificial assumptions, such as an EVM, thus fully leveraging the benefits of data-driven models. Beck et al. [4] suggested that the data-driven SGS model is feasible as long as the direct-closure terms are obtained from a previously conducted DNS. Accordingly, Kurz &\& Beck [26] employed an MLP-based SGS model and showed a good agreement with the fDNS for isotropic turbulence. A similar approach has been adopted by several researchers, such as Gamahara and Hattori [14], Zhou et al. [46], and Liu et al. [29], who predicted $\tau_{ij}$ using resolved flow variables as the input to their MLP-based SGS model. Their results demonstrated that this direct-closure strategy could reproduce the SGS stress distribution in a priori test and turbulence statistics in a posteriori test with a close agreement with the result of the fDNS.
@@
In subsequent studies, instead of applying an MLP algorithm in which both input and output are defined at single-point locations, researchers such as Xie et al. [45] and Park &\& Choi [34] have explored multipoint inputs to improve the predictive accuracy. This approach enables the MLP algorithm to capture turbulent interactions by accounting for non-local input features. Xie et al. [45] used multipoint data of the velocity gradient $\frac{\partial\bar{u}_{i}}{\partial x_{j}}$ as the input and achieved accurate a priori predictions, along with improvements in a posteriori tests when compared with the results of the conventional DSM for isotropic turbulence. Park &\& Choi [34] evaluated both single- and multipoint inputs in a channel flow case. They found that multipoint inputs improved the $\tau_{ij}$ prediction accuracy and achieved a good agreement with the fDNS result in a posteriori test. However, it is worth noting that MLP-based algorithms with multiple inputs often impose a substantial computational burden, making them less effective for large-scale datasets[29].
@@
A more efficient approach to extract multipoint spatial information is through the use of convolutional neural networks (CNN), which capture the spatial correlations via convolutional kernels such that both the memory and computational cost can be reduced, making them more suitable for flow-field data[31, 29]. Given these advantages, CNN-based SGS models have been employed in previous studies, such as those of Beck et al. [4] and Guan et al. [17], for homogeneous isotropic turbulence. Their results demonstrated accurate predictions of $\tau_{ij}$ and showed that the CNN-based models outperformed both static and dynamic Smagorinsky models[4, 17] and the MLP-based SGS models in a posteriori test. Liu et al. [29] employed a CNN-based data-driven SGS model for a turbulent channel flow. They also compared the CNN-based and MLP-based SGS models, showing that the CNN-based model provided more accurate predictions of $\tau_{ij}$. In a posteriori tests, the CNN-based model outperformed the DSM in terms of turbulence statistics. This improvement underscores the advantage of using CNN to capture spatial and turbulent interactions across the neighboring grid points. Furthermore, it is crucial to represent the multiscale characteristics of turbulent flow dynamics[45, 21]. Xie et al.[45] suggested that data-driven SGS models should capture the multiscale characteristic and energy cascade of turbulence through appropriately designed multiple-point input features. To this end, Jalaali and Okabayashi [21] introduced a multiscale CNN-based SGS model designed to capture the vortex interactions across scales, which are essential for energy and momentum transfer. Their model achieved more accurate predictions of $\tau_{ij}$ and improved the turbulent statistics in a posteriori tests when compared with those of both the conventional Smagorinsky model and other CNN-based data-driven SGS models. Moreover, it reproduced the inverse energy cascade and maintained numerical stability in a posteriori tests[21]. Such improvements are likely attributable to the ability of the model to capture interscale vortex interactions, which are not explicitly represented in conventional CNN-based approaches.
@@
Despite these promising developments, for any data-driven SGS model to be physically valid, it must adhere to fundamental invariance principles inherent to turbulence modeling. In modeling terms, the turbulence model must be consistent and frame-indifferent, or satisfy the principle of material objectivity. Fulfilling this requirement has been treated as a hard constraint not only for conventional SGS models[22] but also the data-driven SGS models [42, 13]. As discussed by Oberlack[32], classical SGS models often fail to meet this requirement. In particular, models that explicitly include the rate-of-rotation tensor frequently violate the requirement of frame indifference. To address this, considerable effort has been made to incorporate invariance into the SGS formulations. For example, Oberlack[32] noted that the DSM proposed by Germano et al.[15] satisfied this principle by introducing invariant variables and test-filter quantities that preserved the scaling invariance. Although the model structures considerably differ from those of the conventional SGS models, data-driven SGS models must also adhere to material objectivity, ensuring physically consistent behavior under coordinate rotations. Accordingly, data-driven SGS models are expected to respect the following transformation property:
@@
Likewise, the efforts to satisfy the material objectivity have been undertaken within the data-driven turbulence modeling framework. In this regard, the generalized expansion of the Reynolds stress tensor by Pope [35] has been employed to develop invariant closures for Reynolds stresses. This approach is based on the assumption that the stresses depend on the mean velocity gradient [36]. Within the data-driven framework, this model rewrites the expansion of the Reynolds stress tensor by Pope as
@@
where $T^{(n)}$ are known functions of the symmetric and antisymmetric parts of the velocity gradient tensor, and $\lambda_{1},\ldots,\lambda_{5}$ are scalar invariants. The detailed derivation of these 10 tensors and 5 invariants is given in Pope[36, 35]. The goal is to determine the corresponding coefficients ($c^{(n)}(\lambda_{1},\ldots,\lambda_{5})$), following which the Reynolds stress tensor can be calculated using Eq. (2). The data-driven approach was initially introduced by Ling et al. [27, 28] as the tensor basis neural network (TBNN), and later extended by other researchers, such as Brener et al. [7] in the context of data-driven RANS modeling and Bose &\& Roy[6] for LES closures. These studies relied on scalar invariants to reconstruct the anisotropic stresses and showed a relatively good agreement with the fDNS data.
@@
However, there are limitations to the TBNN approach. As noted by Cinnella [9], the TBNN based on the representation by Pope relies on the hypothesis of the local equilibrium of turbulence, an assumption that no longer holds for flows with strong non-equilibrium. To some extent, the polynomial expansion used in the representation by Pope can be extended to better capture the non-equilibrium turbulence[35, 44]. Prakash et al. [37] noted that the size of the minimal tensor set and the corresponding invariant basis increases with the number of prescribed input features. To address this issue, Prakash et al. [37] proposed a data-driven SGS model based on the eigenstructure of the SGS tensor, using invariant variables as inputs. Their results showed a good agreement with the DNS data for isotropic turbulence. Brener et al. [7] and Wu et al.[44] showed that invariance can also be achieved through data augmentation. By exposing the DNN architecture to multiple orientations of the dataset, the data-driven SGS model was expected to learn representations that are insensitive to frame rotations[44]. However, data augmentation was computationally expensive because it required duplicating the DNS datasets across multiple transformed coordinate systems[7]. Beyond input design and data augmentation, a critical concern remains that invariance should be preserved not only at the level of the input features but also within the structure and operations of the DNN algorithm itself[20, 37, 6]. For example, the study by Jalaali &\& Okabayashi[21] used rotationally invariant inputs, but the architecture of the DNN model itself did not maintain rotational invariance as, a CNN algorithm cannot preserve rotational invariance without the necessary treatment[16]. Because DNN is inherently composed of coordinate-sensitive operations, it is desirable to design architectures or training strategies that are intrinsically invariant.
@@
Motivated by these challenges, we developed a data-driven SGS model that enforces material objectivity and preserves rotational invariance. Here, we adopted the multiscale CNN-based SGS model introduced by Jalaali &\& Okabayashi [21] (MSC model), which captures the flow features across multiple spatial scales. The contributions of this study can be outlined as follows. First, we modified the MSC model by removing the bias terms and batch normalization (BN) layers, following Bin et al. [5] and Cho et al. [8], allowing the model to operate directly on raw input data and enhancing the invariance properties. Second, we integrated a spatial transformer network (STN) [20] into the architecture, enabling the model to learn the features that remain consistent under coordinate rotations. These design choices were evaluated through a priori test with rotated inputs to examine the rotational invariance. The remainder of this paper is organized as follows. In Section II, we introduce the numerical framework and present the proposed data-driven SGS model, with particular emphasis on the overall modeling strategy. Section III describes the results and discussions. Finally, Section IV concludes the study by summarizing the key findings and discussing possible directions for future developments.
@@ II.1 Problem setting
The residual SGS stress, $\tau_{ij}=\overline{u_{i}u_{j}}-\bar{u}_{i}\bar{u}_{j}$, is unclosed and represents the effect of unresolved scales. Therefore, a closure model is required; while by far the most widely used closure adopts the eddy-viscosity hypothesis that can be modeled using conventional approaches such as the Smagorinsky model[40]. In the present study, we replaced such conventional closures with a data-driven SGS model for $\tau_{ij}$. Fig. 1 outlines the modeling workflow. This study was limited to a priori assessment of the capability of the proposed data-driven SGS model to achieve rotational invariance. In a priori tests, the data-driven SGS model is trained in a supervised manner on fDNS flow fields. By using the input data X{X} and label data $\tau_{ij}$, the data-driven SGS model learns a nonlinear mapping such that $\tau_{ij}^{p}=\mathcal{F}({X};w)$, where $\tau_{ij}^{p}$ is the predicted $\tau_{ij}$. The parameter $w$ denotes the weights of the DNN and is obtained by optimization, given as $w=\arg\min_{w}\left(\mathcal{L}(\tau_{ij},\tau_{ij}^{p})\right)$. $\mathcal{L}$ denotes the loss function.